# D4C Glove-train: Solving the RPM and Bongard-logo Problem by Circumscribing and Building Distribution for Concepts

### v9

@@ VIII-A Experiment on RAVEN
Initially, we conducted experiments with Lico-Net on both RAVEN[15] and I-RAVEN[37] datasets to demonstrate its inherent potential. To ensure a fair comparison between excellent models, all experimental settings and equipment in this study were kept consistent with those of RS-Tran [28] and Triple-CFN [32], including data volume, optimizer parameters, data augmentation methods (such as reducing the RAVEN image size to $80\times 80$), batch size, and other relevant hyperparameters.
@@
An interesting observation from the Table II indicates a consensus among existing excellent RPM solvers, including Lico-Net. Specifically, these solvers exhibit differing values of reasoning accuracy when tackling the RAVEN [15] and I-RAVEN [37] datasets. The I-RAVEN dataset was designed to address certain limitations in the option pool configuration of RAVEN by introducing improvements to its negative example setup. These enhancements include, but are not limited to, increasing the number of attributes that are altered in the negative examples and augmenting the magnitude of attribute deviations within them. However, from a human perspective, such additional setups do not alter the difficulty level of reasoning, and thus do not make I-RAVEN any harder or easier compared to RAVEN. This paper speculates that the disparate outcomes achieved by these models on RAVEN and I-RAVEN could potentially be attributed to the superior quality of negative examples present in the I-RAVEN dataset compared to those in RAVEN. Such phenomenons echo the D4C’s spirit.
@@ VIII-B Experiment on Bongard-Logo
We conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN [32] as backbones for D2C, D3C(-cos), and D4C and using PMoC [31] as backbone for D2C and D4C. PMoC, a probabilistic labeling model, does not meet the requirements of D3C(-cos), as D3C(-cos) demands a backbone in the form of $f(z|x)$. The experimental results are presented in the table III.
Upon observing the results presented in the table III, we found that D2C, D3C, and D3C-cos made considerable contributions in the Bongard-Logo task. As can be seen in the table III, replacing D3C with D3C-cos to integrate with ResNet50 would not significantly affect performance. We argue that, from a results-oriented perspective, D3C-cos can be considered as a effective alternative to D3C. Notably, we did not conduct an experiment with Triple-CFN+D3C and Triple-CFN+D2C+D3C. This is because the large number of Triple-CFN parameters makes integrating the D3C approach impractical on our current experimental equipment. This was the driving force that led us to propose D3C-cos as an alternative to D3C in this paper.
@@
Additionally, D4C has imparted varying degrees of improvement to Triple-CFN, ResNet50, and PMoC, with its effectiveness rivaling that of the combined application of D2C and D3C. However, D4C did not yield an ideal improvement on the FF concept for the Bongard-logo problem when applied to Triple-CFN and ResNet50. This study hypothesizes that the suboptimal performance may be attributed to the fact that D4C was specifically designed and tailored for discriminative models with a probabilistic labeling scheme. Conversely, Triple-CFN and ResNet50 do not inherently adhere to such a framework. It has been mentioned that when integrating D4C with Triple-CFN and ResNet50, a decay mechanism was incorporated. This mechanism is a compromise made by D4C to accommodate the non-probabilistic labeling scheme of these models. From the perspective of experimental results, this decay mechanism does not appear to be a fully developed mechanism. Exploring more advanced designs to facilitate the adaptation of D4C to non-probabilistic labeling models, such as Triple-CFN, deserves attention as future work. In contrast to Triple-CFN, PMoC [31] serves as an intuitive probabilistic labeling model for the Bongard-Logo problem. Correspondingly, D4C provides a more significant enhancement when applied to PMoC, owing to its scheme.
D4C demonstrated improvement and contribution, validating the effectiveness of defining tighter problem boundaries for the model with a probabilistic labeling scheme. To integrate D4C with Triple-CFN, we devised a representation decay mechanism that, despite its imperfections, is sufficiently effective. This approach warrants further investigation and development in future research.
Finally, this paper notes that augmentation experiments were conducted for both PMoC [31] and Triple-CFN [32]. These experiments involved expanding the data volume of the Bongard-Logo dataset using data augmentation techniques, specifically including rotation and flipping. It should be emphasized that these data augmentation techniques do not explicitly alter the diversity of concepts or the proportion between concepts inherent in the Bongard-Logo dataset [31]. Rather, the aim of these experiments was to investigate the potential of the models with a larger dataset. Since PMoC posits that Transformer-based visual processing models require substantial data as a foundation, a point also echoed by ViT [17] and RS-Tran [28]. Furthermore, both PMoC and Triple-CFN agree that the utilization of these data augmentation methods serves to enrich the expressions of concepts at the pixel level, thereby challenging the network’s abilities to summarize, derive, accommodate, and invert concepts [31, 32]. This paper conducted experiments to verify the effectiveness of D4C with the expanded dataset. The results are recorded in Table IV.
@@ VIII-C Experiment on PGM
From the experimental results presented in Table V, several notable phenomena emerge. In Triple-CFN, both progressive patterns and minimal reasoning units are positioned at the same logical level [32]. However, a significant design flaw arises from the mismatch in the logical hierarchy between these patterns and units. More specifically, progressive patterns often require the combined expression of multiple minimal reasoning units. Despite this limitation, Triple-CFN still stands as an impressive model. Naturally, leveraging the inductive bias inherent to the RPM problem as the guideline in its design, Lico-Net has achieved even more remarkable results than Triple-CFN in the RPM challenges.
The combination of Lico-Net and D2C, and the combination of Triple-CFN and D2C, did not bring significant improvement, with the enhancement only reflected in the percentile of reasoning accuracy. This phenomenon was anticipated in this paper, as we believe it is due to the difference between RPM and Bongard-Logo problems. In RPM problems, different cases share the same concept space and attribute recognition method. The D2C approach, designed to address issues of concept interference and erroneous coupling among cases, is not applicable to RPM problems.
However, the combination of D4C and D4C glove-train significantly improved Lico-Net’s performance. Additionally, the glove-shaped generator proved to be more suitable for generating negative representations of RPM problems for Lico-Net and enabling it to achieve further refinement. The core innovation of this paper, D4C, has yielded promising results on both the RPM and Bongard-Logo problems, demonstrating the significant contribution of this work.
The combination of Lico-Net and D3C is noteworthy. To validate the theme of D3C, we applied both D3C and D3C-cos to Lico-Net. Through experimentation, we found that D3C and D3C-cos brought improvements to Lico-Net. To demonstrate the effectiveness of D3C’s approach to describing abstract concepts through distributions, this paper conducted ablation experiments using “Lico-Net+D3C-”. Unlike “Lico-Net+D3C”, which encodes the description of image progressive patterns into distributions using a Transformer-Encoder, “Lico-Net+D3C-” encodes them into vectors. To simplify, instead of dividing the encoded representation into mu and logvar, we directly utilize the representation as a vector. Furthermore, the cosine distance between progressive pattern vectors $\{\overline{P}_{nm}\}$ and the vectorized pattern description in Lico-Net is constrained using infoNCE. In essence, the Lico-Net+D3C- can be interpreted as fixing the variance of the distributions established for all concepts in Lico-Net+D3C-cos to 0. The structure of Lico-Net+D3C- is illustrate in figure 19.
Figure 19: Feedforward process of “Lico-Net+D3C-
@@
The D3C’s reasoning accuracys of progressive pattern are recorded in Table VI, highlighting the interpretability expressed by Lico-Net+D3C(-cos). It must be acknowledged that the experiment of applying D3C, which is based on the Sinkhorn distance, to Lico-Net was not conducted to its fullest extent in this paper. Once the experimental results of Lico-Net+D3C were sufficient to illustrate our point, we halted the model training. This was due to the significant computational resources required, even with the experimental equipment used in this paper: four A100s graphics cards. The training involved numerous epochs and tedious individual training epochs. This was the intended purpose behind the design of D3C-cos as a replacement for D3C, as the power of Zeus needs to be conveyed by Hermes.
Despite achieving remarkable results in terms of reasoning accuracy and interpretability of progressive patterns with Lico-Net+D3C or D3C-cos, it is noteworthy that the improvement over previous works involving Meta data as additional auxiliary labels is not particularly striking. The possible reason for this is that the introduction of Meta data significantly reduces the difficulty of the RPM problem Moreover, it is anticipated that a greater abundance of supervisory signals would correlate with enhanced intelligent performance.
@@
However, Lico-Net+D3C still demonstrates distinct advantages that surpass other methods involving metadata. Specifically, Tranclip [28] utilizes metadata to pre-train a set of warm-start parameters for RS-Tran [28]. In other words, Tranclip fails to simultaneously endow RS-Tran with both interpretability of progressive patterns and high reasoning accuracy. Meta Triple-CFN [32] also relies on its specially designed Re-space Layer to enhance both its interpretability and reasoning accuracy. Similarly, the Re-Space Layer requires a fully trained Meta Triple-CFN to provide warm-start parameters [32]. In contrast, Lico-Net operates effectively without the necessity of a Re-space layer or a warm-start process, thereby simplifying the training procedure to a certain extent. This indicates that Lico-Net and D3C are more direct and effective, eliminating the need for indirect approaches. This ”directness” reflects a more coherent and self-consistent logic of Lico-Net and D3C compared to other methods incorporating metadata.
This paper argues that this is not an insignificant contribution, as warm-start behavior may introduce unexpected instabilities, which are not a concern for Lico-Net and D3C. For example, when Tranclip [28] uses Meta data to produce warm-start parameters for RS-Tran, early stopping is required [28]. Neglecting this step may hinder Tranclip from improving or potentially worsen the reasoning performance of RS-Tran. The Rs-space Layer [32] relies on a well-trained Meta Triple-CFN to provide warm-start parameters [32], a process that demands considerable training time and epochs. Inadequate training of these warm-start parameters can expose Meta Triple-CFN to risks akin to those faced by RS-Tran. The stopping timing for the preparation of both warm-start parameters requires manual determination and control, which brings numerous potential issues. In contrast, Lico-Net effectively mitigates these potential issues.
VIII-D Ablation experiment on D2C and D4C
In the final of this section, we devised an experiment to validate the D2C and D4C approach. Utilizing a 500-step DDPM model, we generated a substantial quantity of Bongard-Logo data, specifically 3,180 Bongard-Logo images. According to the DDPM theory, these images belong to the same distribution as the real data, albeit with unknown semantics and concepts. During the training of ResNet50 and Triple-CFN, for each feedforward pass, we introduced 14 randomly selected DDPM-generated images as auxiliary group examples for the Bongard-Logo cases. The experimental results were recorded in “D2C-” item of the Table VII.
TABLE VII: Reasoning Accuracies of models on Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
ResNet50 (infoNCE baseline) 92.0 88.9 98.0 76.1 76.0
Triple-CFN 93.2 92.0 98.2 78.0 78.1
ResNet50+D2C- 91.6 90.0 97.9 76.3 76.1
Triple-CFN+D2C- 93.0 92.1 98.2 77.9 78.2
It can be observed from the experiment results recorded in Table VII that introducing generated samples as negative examples is difficult to fundamentally alter the network’s performance, as such behavior struggles to re-delineate the concepts. DDPM-generated data, sharing the same distribution as real data but with inherent uncertainties of concept and attribution, are limited in contributing to the delineation of concept boundaries.
D4C posits that the accuracy achieved by the discriminator in abstract reasoning problems is closely related to the quality of the positive and negative examples set in the task. The established examples serve to delineate the concept boundaries for abstract reasoning challenges. D2C, while attempting to define new concept boundaries for such problems, represents an imperfect methodology. Conversely, D4C emerges as a relatively more comprehensive approach. This ablation experiment aims to verify that, within the context of Bongard-Logo problems, the provision of meaningless negative examples—examples that do not meaningfully contribute to defining the boundaries of reasoning tasks—fails to enhance the discriminator’s effectiveness. This finding echoes similar observations made in the RAVEN dataset, reinforcing the significance and rationale behind the design of D4C. Furthermore, it reaffirms D4C’s assertion regarding the importance of high-quality positive and negative examples in abstract reasoning tasks. It also provides indirect evidence that the approaches of D2C and D4C extend beyond merely introducing additional negative examples. Instead, they aim to introduce negative examples that better delineate the boundaries of the problem.
@@ IX Conclusion
In summary, D3C advocates for representing concepts in abstract reasoning problems using distributions. It posits that, compared to vectors, distributions are better suited to capture the high-dimensionality, abstractness, and complexity inherent in human concepts, and experimental evidence underscores the superiority of this approach. Additionally, D2C presents a novel approach to tackling Bongard-Logo and RPM problems by resampling training data and introducing new loss function terms, effectively outlining new boundaries for solutions to these challenges. However, the extent to which these boundaries adapt to the specific problems remains uncertain.
D4C builds upon the principles of both D2C and D3C, exploring the roles of positive and negative examples in abstract reasoning problems from a perspective on distribution. Specifically, positive examples are utilized to depict the mean for deep networks, while negative examples contribute to describing the variance. Based on this understanding, D4C employs an adversarial approach, pitting the generator against the discriminator in repeated challenges. The generator continually produces negative examples that push the boundaries of the problem, while the discriminator refines its skills through repeated defense. This iterative process not only enhances the discriminator’s ability to distinguish between positive and negative examples but also improves the generator’s capacity to generate more sophisticated negative examples. Ultimately, this adversarial framework promotes the evolution of both components, leading to more robust and reliable solutions in abstract reasoning tasks.
For future, the D4C method presents a novel approach for abstract reasoning problems and even other discriminative tasks. Its exploration of the identities of positive and negative examples may provide insights for the development of other works.

### v12

@@ VIII-A Experiment on RAVEN
Initially, we tested Lico-Net on RAVEN and I-RAVEN datasets to show its potential. For a fair comparison, we used the same settings and equipment as RS-Tran[28] and Triple-CFN[32], including data volume, optimizer parameters, data augmentation, batch size, and other hyperparameters.
@@
An interesting observation from the Table II indicates a commonality among existing excellent RPM solvers, including Lico-Net. Specifically, these solvers exhibit varying reasoning accuracy when tackling the RAVEN [15] and I-RAVEN [37] datasets. The I-RAVEN dataset was designed to address certain limitations in the option pool configuration of RAVEN by introducing improvements to its negative example setup. These enhancements include, but are not limited to, increasing the number of attributes that are altered in the negative examples and augmenting the magnitude of attribute deviations within them. However, from a human perspective, since the setup of positive examples remains unchanged, such additional setups do not alter the difficulty of reasoning, making I-RAVEN neither more nor less challenging than RAVEN. This paper speculates that the disparate outcomes achieved by these models on RAVEN and I-RAVEN could potentially be attributed to the higher quality of negative examples present in the I-RAVEN dataset compared to those in RAVEN. Such phenomena echo the D4C’s spirit.
@@ VIII-B Experiment on Bongard-Logo
We conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN [32] as backbone for D2C, D3C(-cos), and D4C and using PMoC [31] as backbone for D2C and D4C. PMoC, a probabilistic labeling model, does not meet the requirements of D3C(-cos), as D3C(-cos) demands a backbone in the form of $f(z|x)$. The experimental results are presented in the table III.
Observing table III, D2C and D3C(-cos) notably contributed to the Bongard-Logo task.Replacing D3C with D3C-cos when integrated with ResNet50 has minimal impact on performance. From the result-oriented perspective, D3C-cos is a qualified alternative for D3C. We didn’t test Triple-CFN with D3C or D2C+D3C due to practical limitations from our equipment, which led to the proposal of D3C-cos. Furthermore, D4C has significantly improved Triple-CFN, ResNet50, and PMoC, often matching the combined effects of D2C and D3C. However, on Bongard-Logo’s FF concept, D4C’s improvements on Triple-CFN and ResNet50 were less than ideal. This might be because D4C was designed for probabilistic labeling models, whereas Triple-CFN and ResNet50 do not follow this probabilistic approach. A decay mechanism, as a compromise for non-probabilistic labeling models, was added when integrating D4C with these models. Experimental results suggest this mechanism needs further refinement. Future work should explore better designs to adapt D4C to non-probabilistic labeling models like Triple-CFN. Conversely, PMoC [31], a probabilistic labeling model for Bongard-Logo, benefits more from D4C due to its scheme alignment.
@@
Finally, this paper notes that augmentation experiments were conducted for both PMoC [31] and Triple-CFN [32]. These experiments involved expanding the data volume of the Bongard-Logo dataset using data augmentation techniques, specifically including rotation and flipping. It should be emphasized that these augmentation techniques do not explicitly alter the diversity of concepts or the proportion between concepts inherent in the Bongard-Logo dataset [31]. Rather, the aim of these experiments was to investigate the potential of the models with a larger dataset. Since PMoC posits that Transformer-based visual processing models require substantial data as a foundation, a point that is also echoed by ViT [17] and RS-Tran [28]. Furthermore, both PMoC and Triple-CFN agree that the utilization of these data augmentation methods serves to enrich the expressions of concepts at the pixel level, thereby challenging the network’s abilities to summarize, derive, accommodate, and invert concepts [31, 32]. To verify the effectiveness of D4C on the augmented dataset, we conducted experiments, and the results are presented in Table IV.
@@ VIII-C Experiment on PGM
From the experimental results presented in Table V, several notable phenomena emerge. Triple-CFN[32] positions progressive patterns and minimal reasoning units at the same logical level, leading to a design flaw due to their mismatched logic. Despite this, Triple-CFN remains an impressive model. By leveraging the RPM problem’s accurate inductive bias in its design, Lico-Net has surpassed Triple-CFN in RPM challenges.
The combinations of Lico-Net+D2C and Triple-CFN+D2C showed minimal improvement, limited to reasoning accuracy percentile. This was expected, as RPM and Bongard-Logo differ. RPM cases share a concept space and attribute recognition, rendering D2C, designed for concept interference and erroneous coupling issues, ineffective for RPM. The combination of D4C and D4C glove-train notably enhanced Lico-Net’s performance. D4C, our core contribution, showed promising results on both RPM and Bongard-Logo problems, highlighting the significance of this work.
The combination of Lico-Net and D3C is noteworthy. To validate D3C method, we applied both D3C and D3C-cos to Lico-Net. Through experimentation, we found that D3C and D3C-cos brought improvements to Lico-Net. To demonstrate the effectiveness of D3C’s approach to describing abstract concepts through distributions, this paper conducted ablation experiments using “Lico-Net+D3C-”. Unlike “Lico-Net+D3C”, “Lico-Net+D3C-” encodes Meta data into vectors. In essence, the Lico-Net+D3C- can be interpreted as fixing the variance of the distributions established for all concepts in Lico-Net+D3C-cos to 0.
@@
Table VI records D3C’s reasoning accuracies for progressive patterns, emphasizing Lico-Net+D3C(-cos)’s interpretability. Due to resource constraints, this paper did not fully explore the application of Sinkhorn-based D3C to Lico-Net. We stopped training once results illustrated our point, given the significant computational demands even with four A100s. In terms of results, D3C-cos can serve as a stand of D3C. Despite impressive reasoning accuracy and interpretability with Lico-Net+D3C(-cos), the improvement over previous works using Meta data as auxiliary labels isn’t significant. This might be because the inclusion of Meta data lowers the RPM problem’s reasoning difficulty. More supervisory signals typically enhance the intelligent performance of the system.
@@
Lico-Net+D3C still shows distinct advantages over other Meta data-based methods. Specifically, Tranclip[28] utilizes Meta data to pre-train a set of warm-start parameters for RS-Tran[28], achieving promising reasoning accuracy for RS-Tran. However, Tranclip fails to simultaneously endow RS-Tran with both the the interpretability of progressive patterns and high reasoning accuracy. Meta Triple-CFN[32] also relies on a specially designed Re-space Layer[32] to enhance both its interpretability and reasoning accuracy. Similarly, the Re-Space Layer requires a fully trained Meta Triple-CFN to provide warm-start parameters [32]. In contrast, Lico-Net+D3C operates effectively without these extras, showing a more direct, coherent and self-consistent logic. This paper argues that this is not an insignificant contribution, as warm-start behavior may introduce unexpected instabilities, which are not a concern for Lico-Net and D3C. For instance, Tranclip requires early stopping when using Meta data for training RS-Tran’s warm-start parameters[28], and neglecting this can hinder RS-Tran’s performance. Similarly, The Re-space Layer [32] relies on a well-trained Meta Triple-CFN to provide warm-start parameters [32], a process that demands considerable training time and epochs, with inadequate training posing risks. Both methods require manual control, whereas Lico-Net+D3C mitigates these issues.
@@ IX Conclusion
D4C, building on D2C and D3C principles, explores the roles of positive and negative examples in abstract reasoning from a distribution perspective. Positive examples depict the mean for solution distribution, while negative ones limit variance. This seems to reflect a despairing fact: the upper limit of accuracy achieved by the discriminator may be constrained by the quality of positive and negative examples. Therefore, D4C emerged to address this phenomenon. D4C adopts an adversarial approach, where the generator creates challenging negative examples and the discriminator refines its skills through defense. This iterative process enhances both components, resulting in remarkable solutions for abstract reasoning tasks. This paper does not advocate for the interpretability of the samples generated by the D4C generator; they are solely used to improve the performance of the discriminator. For future, the D4C method presents a unique perspective and approach for abstract reasoning problems and even other discriminative tasks. Its exploration of the identities of positive and negative examples may provide insights for the development of other works.